# Simple and Effective Specialized Representations for Fair Classifiers

### v1

@@ 1 Introduction
To better balance fairness and predictive accuracy, many approaches aim to learn fair representations by projecting input data into a latent space that obscures sensitive information while preserving task-relevant structure zemel2013learning ; madras2018learning ; moyer2018invariant ; balunovic2022fair . These approaches typically fall into two categories, which we refer to as specialized and general. Specialized representations are tailored to a specific task, filtering out sensitive and irrelevant features for improved performance. General representations, in contrast, are task-agnostic and aim for broad applicability across multiple tasks while being insensitive to protected attributes, but are harder to learn and may compromise predictive accuracy.
@@ 3 Background
Since strict fairness criteria often entail accuracy trade-offs due to correlation between the sensitive $S$ and the task label $Y$, fair representations typically result in either performance reductions or leakage of the sensitive information in the classification. Consequently, we adopt the concept of $\epsilon$-fairness, where $\epsilon$ represents the statistical distance between the learned representations $\mathbb{P}_{Z\mid S}$. This enables the application of bounds from Madras et al. madras2018learning relating statistical distances to several fairness metrics (see section A.4). Specifically, they consider the Total Variation (TV) distance, defined for distributions $\mathbb{P}_{0}$ and $\mathbb{P}_{1}$ as:
[…]
@@
$\sup_{\mathcal{F}}\max_{f\in\mathcal{F}}\mathbb{P}_{(}Y=f(X))=\frac{1}{2}\left(1+\Delta(\mathbb{P}_{Z\mid s=0},\mathbb{P}_{Z\mid s=1})\right).$ (2)
Adversarial Evaluation should not be confused with Adversarial Learning. Adversarial Learning learns a latent fair representation $Z$ from $X$ optimizing eq. 2 directly via a min-max training, which can lead to training instability feng2019learning ; moyer2018invariant ; elazar2018adversarial . Adversarial frameworks are typically robust only to the class of functions used during training. Adversarial Evaluation simply assesses the fairness of a fixed representation $Z$ by quantifying how accurately a classifier (often an Multi Layer Perceptron (MLP)) can predict the sensitive attribute $S$ from $Z$.
@@ 4 Fair Representations matching Characteristic Function
As discussed in section 3, to incentivize fair representation, we need to match the different conditional distributions $\mathbb{P}_{Z\mid S}$, thus minimizing their relative statistical distance. To achieve this, we propose the addition of a differentiable penalty term based on the Characteristic Function Distance. We refer to the overall approach as Fairness matching Characteristic Function (FmCF).
@@
[…]
The CF is closely related to the notion of Fourier Transform, and inherits several important properties. It always exists, it’s bounded $|\varphi_{X}(t)|\leq 1$, but most importantly, there is a one-to-one correspondence between probability measures and CFs lukacs1970characteristic . That is, given two random variables $X$ and $Y$, with probability measures $\mathbb{P}_{X}$ and $\mathbb{P}_{Y}$ then it holds that:
@@
[…]
@@ 5 Fair Classification matching Sufficient Statistics
The method introduced in section 4 provides a versatile framework for learning specialized representations $Z$ from $X$ independent from $S$ for a general task. However, in practice, most fairness-aware scenarios involve classification wu2018fairness ; madhavan2020fairness ; zafar2017fairness . By simplifying the approach, tailoring it to fair classification, we can provide provable post-hoc fairness guarantees. Furthermore, we can also make it computationally cheaper, relaxing the need for a Monte Carlo estimate of the CF.
@@
The penalty term introduced in section 4 involves sampling multiple points $t\sim\mathbb{P}_{T}$ from the weighting kernel where the empirical CF is evaluated at. The number of points needed for a reliable estimate of the CFD grows rapidly in the number of dimensions ansari2020characteristic .
@@
Assuming that $\varphi_{X}$ is analytic, matching the derivatives evaluated at $t=0$ would be sufficient to ensure fairness, without need for sampling from $\mathbb{P}_{T}$. There exist statistical tools to estimate the full series (section A.3); however, they might lead to strong instabilities in training.
A pragmatic approach is to truncate the expansion to some order $N$. This objective does not guarantee that the two distributions are exactly equal, however, we show that truncating to $N=2$ is sufficient to ensure fair classification.
@@
A generic MLP with $M$ layers used for classification can be interpreted as $M-1$ layers of encoder $z=h_{\theta}(x)$, and a final layer of a LR $\hat{y}=\sigma_{\beta}(z)$. Therefore, by applying a fairness penalty to $z$, we can restrict the analysis to the family of LR classifiers.
@@
Algorithm 1 FmCF loss term
1:  Input: encoder $h_{\theta}$, decoder $f_{\theta}$, batch =$\mathcal{B}=\{(x_{i},y_{i},s_{i})\sim\mathbb{P}_{X,Y,S}\}$.
2:  $z\leftarrow h_{\theta}(x)$
3:  $\hat{y}\leftarrow f_{\theta}(z)$
4:  for all $s\in S$ do
5:     for all $j\in\{1,\dots,k\}$ do
6:        Sample $t_{j}\sim\mathbb{P}_{T}$
9:     end for
10:  end for
11:  $\mathcal{L}_{\text{CF}}\leftarrow\displaystyle\sum_{s\in\mathcal{S}}\frac{1}{k}\sum_{j=1}^{k}\left\lvert\varphi_{\mathcal{N}}(t_{j})-\hat{\varphi}_{Z\mid s}(t_{j})\right\rvert^{2}$
12:  $\mathcal{L}\leftarrow\mathcal{L}_{C}(\hat{y},y)+\alpha\,\mathcal{L}_{\text{CF}}$
Algorithm 2 FmSS Training
1:  Input: encoder $h_{\theta}$, linear regression $f_{\beta}$, batch =$\mathcal{B}=\{(x_{i},y_{i},s_{i})\sim\mathbb{P}_{X,Y,S}\}$.
2:  $z\leftarrow h_{\theta}(x)$
3:  $\hat{y}\leftarrow f_{\beta}(z)$
4:  for all $s\in S$ do
5:     $\sigma_{s}^{2}\leftarrow\operatorname{Var}[Z\mid S=s]$
6:     $\mu_{s}\leftarrow\mathbb{E}[Z\mid S=s]$
7:  end for
8:  $\mathcal{L}_{\text{KL}}\leftarrow\displaystyle\sum_{s\in\mathcal{S}}||\sigma_{s}^{2}+\mu_{s}^{2}-1-\log\sigma_{s}^{2}||_{1}$
9:  $\mathcal{L}\leftarrow\mathcal{L}_{C}(\hat{y},y)+\alpha\,\mathcal{L}_{\text{KL}}$
@@
Theorem 5.2 shows that, building an intermediate representation $Z$ with $\mathbb{E}_{Z\mid S}\left[Z\right]=\mathbb{E}_{Z}\left[Z\right]$ is sufficient to ensure that LR models do not use information about the sensitive attribute $S$.
However, we can derive an even stronger guarantee, analyzing what happens for ${\mathbb{E}_{Z}\left[Z_{\left[i\right]}\mid S\right]\approx\mathbb{E}_{Z}\left[Z_{\left[i\right]}\right]}$. In particular, the mean and variance of $\mathbb{P}_{Z\mid S}$ are enough to fully characterize the predictive power of a feature.
@@
One empirical example of the consequences of theorem 5.3 is shown in fig. 2 where a Gaussian and a mixture of Gaussians are used to model $\mathbb{P}_{Z\mid S}$, and $\mathbb{E}_{Z\mid S}\left[Z\right]$ gradually converge to the same value, while the variance is preserved.
Similarly to section 4 we adopt a Gaussian target distribution for $\mathbb{P}_{Z\mid S}$ This is a particularly compelling choice from the lens of the maximum entropy principle [^1] When constrained by fixed first and second moments, the Gaussian distribution uniquely maximizes entropy..
@@
This approach fundamentally departs from popular methods in the literature, which primarily focus on minimizing $\Delta(\mathbb{P}_{Z\mid S=0},\mathbb{P}_{Z\mid S=1})$, by permitting the sensitive attribute $S$ to be encoded within the learned representation. Crucially, it ensures that $S$ is embedded in a manner that renders it provably inaccessible to the classifier being trained.
Leveraging the convexity of LR, one can precisely quantify the $\epsilon$-fairness of a given representation $z$, thus establishing that no logistic classifier $\sigma_{\beta}(x)$ can extract more than an $\epsilon$ amount of information regarding the sensitive attribute $s$. This constitutes a substantially stronger guarantee than adversarial evaluation, which lacks provable tightness and robustness.
@@ 6 Experimental Evaluation
FNF balunovic2022fair offers the strongest theoretical and empirical guarantees, making it the current state of the art in fair representation learning. Therefore, we primarily evaluate the performance of our proposed methods on the same suite of benchmarks, while also considering additional methods and datasets for a more comprehensive analysis. In all tested datasets, both proposed approaches match or surpass the performances of state-of-the-art approaches in terms of accuracy, while providing fairer representations. Surprisingly, even though the approach presented in section 5 gives guarantees against the worst-case linear attacker, it still holds good performances against adversarial evaluations using deep neural networks. Other results on different setups can be found in section A.5, where we additionally report more Pareto plots using different hyperparameters for both FmSS (section 5) and FmCF (section 4)
@@
This dataset is the most fundamental benchmark in fairness, as almost all methods in literature are evaluated on it jaiswal2020invariant ; roy2019mitigating ; madras2018learning ; balunovic2022fair . It requires predicting the income of a person without considering their gender ding2021retiring . For this dataset, we use the same setup as balunovic2022fair , by evaluating our approach both on a classifier in the same family of the classifier used for the task, and one from a family of much more expressive ones, thus deeper and wider. This allows for a much tighter evaluation of the statistical distance $\Delta(\mathbb{P}_{Z\mid S=0},\mathbb{P}_{Z\mid S=1})$. However, in section A.5 we compare to other baselines that only offer guarantees on the trained family of functions. For the proposed methods in table 1 $g\in\mathcal{G}$ is the class of LR classifiers, while $g\not\in\mathcal{G}$ are deep MLPs (see section A.7 for details). By construction, FmCF should be evaluated for $g\not\in\mathcal{G}$, while FmSS for $g\in\mathcal{G}$.
@@
The German dataset has been explored in the literature under different settings. In particular, moyer2018invariant considers gender as the sensitive attribute, with the task of predicting income based on a set of personal features. To ensure a fair comparison, we evaluate our approach against several existing methods using the same experimental setup as in moyer2018invariant , as presented in table 2. The proposed method outperforms the other approaches both in terms of accuracy and fairness, showcasing the effectiveness of the novel penalties. Further comparison under sensitive attributes is presented in section A.5.
@@
Since balunovic2022fair achieves the best balance between fairness and accuracy, in fig. 3 we compare our proposed method from section 4 against all datasets used in their evaluation. These include four distinct datasets: Crime and Health, which contain continuous features that NFs can handle directly, and Compas and Adult, which include discrete features requiring separate handling by Normalizing Flows.
The plots demonstrate that our proposed method consistently outperforms or matches balunovic2022fair , with the added advantage of not requiring the sensitive attribute during evaluation. Additionally, in section A.5, we report the Pareto fronts obtained by varying the latent dimension size, highlighting the impact on performance.
@@
In order to have a better sense of what the CF approach presented in section 4 is learning, we visualize the learned representation in fig. 4.
@@ 7 Conclusions
In this work, we introduce a novel framework for learning fair specialized representations using CFD, addressing key limitations of existing adversarial and normalizing flow-based approaches. By leveraging CFs, our method ensures stable and efficient fairness by minimizing sensitive information leakage while maintaining high predictive performance. We also present a simplified version of the framework that allows for certifiable bounds on the amount of sensitive information used by downstream tasks. Experimental results show that our approaches consistently outperform or match state-of-the-art methods in terms of the fairness-accuracy trade-off across a range of benchmark datasets. Unlike many existing approaches, our method does not require access to sensitive attributes at inference time, making it more practical for real-world applications.
Our approach opens a new avenue for research in fair representation learning by offering a principled, non-adversarial alternative to existing methods. However, like many fairness-focused algorithms, it faces challenges in high-dimensional settings, where the CFD can become less effective. This limitation underscores the need for further research to improve the scalability and robustness of our method in such scenarios. Additionally, current fairness benchmarks largely focus on low-dimensional datasets, limiting the ability to comprehensively evaluate and compare methods in more realistic and complex settings. Developing more sophisticated, high-dimensional fairness datasets would greatly enhance the evaluation of approaches like ours and drive further innovation in the field.
8 Broader Impact

### v2

@@ 1 Introduction
To better balance fairness and predictive accuracy, many approaches aim to learn fair representations by projecting input data into a latent space that obscures sensitive information while preserving task-relevant structure zemel2013learning ; madras2018learning ; moyer2018invariant ; balunovic2022fair . Once optimized, these representations can replace the original data during both training and testing, enabling fair decision-making in scenarios where sensitive attributes must not be explicitly used, such as determining loan approvals or hiring decisions. These approaches typically fall into two categories, which we refer to as specialized and general. Specialized representations are tailored to a specific task, filtering out sensitive and irrelevant features for improved performance. General representations, in contrast, are task-agnostic and aim for broad applicability across multiple tasks while being insensitive to protected attributes, but are harder to learn and may compromise predictive accuracy. Returning to the loan approval example, specialized representations learned in that context would be suitable only for predicting creditworthiness. If reused for another task, e.g. default risk or customer lifetime value, they might lack task-relevant information that was discarded during the fairness optimization. General representations, instead, preserve information that remains broadly useful across prediction tasks while still mitigating sensitivity to protected attributes.
@@ 3 Background
Since strict fairness criteria often entail accuracy trade-offs due to correlation between the sensitive $S$ and the task label $Y$, fair representations typically result in either performance reductions or leakage of the sensitive information in the classification. Consequently, we adopt the concept of $\epsilon$-fairness, where $\epsilon$ represents the statistical distance between the learned representations $\mathbb{P}_{Z\mid S}$. This enables the application of bounds from Madras et al. madras2018learning relating statistical distances to several fairness metrics (see section˜A.4). Specifically, they consider the Total Variation (TV) distance, defined for distributions $\mathbb{P}_{0}$ and $\mathbb{P}_{1}$ as:
$\Delta(\mathbb{P}_{0},\mathbb{P}_{1})=\frac{1}{2}\int_{\mathcal{X}}|\mathbb{P}_{0}(x)-\mathbb{P}_{1}(x)|\,dx.$ (1)
@@
$\sup_{\mathcal{G}}\max_{g\in\mathcal{G}}\mathbb{P}(Y=g(X))=\frac{1}{2}\left(1+\Delta(\mathbb{P}_{Z\mid s=0},\mathbb{P}_{Z\mid s=1})\right).$ (2)
Adversarial Evaluation should not be confused with Adversarial Learning. Adversarial Learning learns a latent fair representation $Z$ from $X$ optimizing eq.˜2 directly via a min-max training, which can lead to training instability feng2019learning ; moyer2018invariant ; elazar2018adversarial . Adversarial frameworks are typically robust only to the class of functions used during training. Adversarial Evaluation simply assesses the fairness of a fixed representation $Z$ by quantifying how accurately a classifier (often an Multi Layer Perceptron (MLP)) can predict the sensitive attribute $S$ from $Z$.
@@ 4 Fair Representations matching Characteristic Function
Algorithm 1 FmCF loss term
1:Input: encoder $h_{\theta}$, predictor $f_{\theta}$, batch $\mathcal{B}=\{(x_{i},y_{i},s_{i})\sim\mathbb{P}_{X,Y,S}\}$.
[…]
4:for all $s\in S$ do
5:  for all $j\in\{1,\dots,k\}$ do
6:   Sample $t_{j}\sim\mathbb{P}_{T}$
9:  end for
10:end for
[…]
The concept of Characteristic Function has been extensively studied for a long time in the statistical testing literature lukacs1970characteristic , but interest in its application to ML has only recently grown ansari2020characteristic . As discussed in section˜3, to incentivize fair representation, we need to match the different conditional distributions $\mathbb{P}_{Z\mid S}$, thus minimizing their relative statistical distance. To achieve this, we propose the addition of a differentiable penalty term based on the Characteristic Function Distance. We refer to the overall approach as Fairness matching Characteristic Function (FmCF).
@@
$\varphi_{X}(t)=\mathbb{E}_{X}\left[e^{i\langle X,t\rangle}\right]=\int_{x\in\mathbb{R}^{d}}e^{i\langle x,t\rangle}d\,\mathbb{P}_{X}.$ (5)
The CF is closely related to the notion of the Fourier Transform, and inherits several important properties. It always exists, it’s bounded $|\varphi_{X}(t)|\leq 1$, but most importantly, there is a one-to-one correspondence between probability measures and CFs lukacs1970characteristic . That is, given two random variables $X$ and $Y$, with probability measures $\mathbb{P}_{X}$ and $\mathbb{P}_{Y}$ then it holds that:
@@
$\varphi_{\mathcal{N}}(t)=\int_{x\in\mathbb{R}^{d}}(2\pi)^{\frac{-d}{2}}e^{-\frac{\left\lVert x\right\rVert^{2}}{2}}e^{i\langle x,t\rangle}dx=e^{-\frac{\left\lVert t\right\rVert^{2}}{2}}.$ (8)
@@ 5 Fair Classification matching Sufficient Statistics
Algorithm 2 FmSS Training
1:Input: encoder $h_{\theta}$, linear regression $f_{\beta}$, batch $\mathcal{B}=\{(x_{i},y_{i},s_{i})\sim\mathbb{P}_{X,Y,S}\}$.
[…]
4:for all $s\in S$ do
5:  $\sigma_{s}^{2}\leftarrow\operatorname{Var}[Z\mid S=s]$
6:  $\mu_{s}\leftarrow\mathbb{E}[Z\mid S=s]$
7:end for
[…]
The method introduced in section˜4 provides a versatile framework for learning specialized representations $Z$ from $X$ independent from $S$ for a general task. However, in practice, most fairness-aware scenarios involve classification wu2018fairness ; madhavan2020fairness ; zafar2017fairness . By simplifying the approach, tailoring it to fair classification, we can provide provable post-hoc fairness guarantees. Furthermore, we can also make it computationally cheaper, relaxing the need for a Monte Carlo estimate of the CF.
@@
The penalty term introduced in section˜4 involves sampling multiple points $t\sim\mathbb{P}_{T}$ from the weighting kernel where the empirical CF is evaluated at. The number of points needed for a reliable estimate of the CFD grows rapidly in the number of dimensions ansari2020characteristic .
@@
Assuming that $\varphi_{X}$ is analytic, matching the derivatives evaluated at $t=0$ would be sufficient to ensure fairness, without the need for sampling from $\mathbb{P}_{T}$. There exist statistical tools to estimate the full series (section˜A.3); however, they might lead to strong instabilities in training.
A pragmatic approach is to truncate the expansion to some order $N$. This objective does not guarantee that the two distributions are exactly equal; however, we will show that truncating to $N=2$ is sufficient to ensure fair classification.
@@
A generic MLP with $M$ layers used for classification can be interpreted as M1M-1 layers of encoder $z=h_{\theta}(x)$, and a final layer of a LR $\hat{y}=\sigma_{\beta}(z)$. Therefore, by applying a fairness penalty to $z$, we can restrict the analysis to the family of LR classifiers.
@@
Theorem˜5.2 shows that, building an intermediate representation $Z$ with $\mathbb{E}_{Z\mid S}\left[Z\right]=\mathbb{E}_{Z}\left[Z\right]$ is sufficient to ensure that LR models do not use information about the sensitive attribute $S$.
However, it is possible to derive an even stronger guarantee, analyzing what happens for $\mathbb{E}_{Z}\left[Z_{\left[i\right]}\mid S\right]\approx\mathbb{E}_{Z}\left[Z_{\left[i\right]}\right]$}. In particular, theorem˜5.3 highlights that the mean and variance of $\mathbb{P}_{Z\mid S}$ are enough to fully characterize the predictive power of a feature.
@@
One empirical example of the consequences of theorem˜5.3 is shown in fig.˜2 where a Gaussian and a mixture of Gaussians are used to model $\mathbb{P}_{Z\mid S}$, and $\mathbb{E}_{Z\mid S}\left[Z\right]$ gradually converge to the same value, while the variance is preserved.
Similarly to section˜4 we adopt a Gaussian target distribution for $\mathbb{P}_{Z\mid S}$ This is a particularly compelling choice from the lens of the maximum entropy principle [^1] When constrained by fixed first and second moments, the Gaussian distribution uniquely maximizes entropy..
@@
This approach fundamentally departs from popular methods in the literature, which primarily focus on minimizing $\Delta(\mathbb{P}_{Z\mid S=0},\mathbb{P}_{Z\mid S=1})$, by permitting the sensitive attribute $S$ to be encoded within the learned representation. Crucially, it ensures that $S$ is embedded in a manner that renders it provably inaccessible to the classifier being trained. Leveraging the convexity of LR, one can precisely quantify the $\epsilon$-fairness of a given representation $z$, thus establishing that no logistic classifier $\sigma_{\beta}(x)$ can extract more than an $\epsilon$ amount of information regarding the sensitive attribute $s$. This constitutes a substantially stronger guarantee than adversarial evaluation, which lacks provable tightness and robustness.
@@ 6 Experimental Evaluation
FNF balunovic2022fair offers the strongest theoretical and empirical guarantees, making it the current state of the art in fair representation learning. Therefore, we primarily evaluate the performance of our proposed methods on the same suite of benchmarks, while also considering additional methods and datasets for a more comprehensive analysis. In all tested datasets, both proposed approaches match or surpass the performances of state-of-the-art approaches in terms of accuracy, while providing fairer representations. Surprisingly, even though the approach presented in section˜5 gives guarantees against the worst-case linear attacker, it still holds good performances against adversarial evaluations using deep neural networks. Other results on different setups can be found in section˜A.5, where we additionally report more Pareto plots using different hyperparameters for both FmSS (section˜5) and FmCF (section˜4)
@@
This dataset is the most fundamental benchmark in fairness, as almost all methods in literature are evaluated on it jaiswal2020invariant ; roy2019mitigating ; madras2018learning ; balunovic2022fair . It requires predicting the income of a person without considering their gender ding2021retiring . For this dataset, we use the same setup as balunovic2022fair , by evaluating our approach both on a classifier in the same family as the classifier used for the task, and one from a family of much more expressive ones, thus deeper and wider. This allows for a much tighter evaluation of the statistical distance $\Delta(\mathbb{P}_{Z\mid S=0},\mathbb{P}_{Z\mid S=1})$. However, in section˜A.5 we compare to other baselines that only offer guarantees on the trained family of functions. For the proposed methods in table˜1 $g\in\mathcal{G}$ is the class of LR classifiers, while $g\not\in\mathcal{G}$ are deep MLPs (see section˜A.7 for details). By construction, FmCF should be evaluated for $g\not\in\mathcal{G}$, while FmSS for $g\in\mathcal{G}$.
@@
The German dataset has been explored in the literature under different settings. In particular, moyer2018invariant considers gender as the sensitive attribute, with the task of predicting income based on a set of personal features. To ensure a fair comparison, we evaluate our approach against several existing methods using the same experimental setup as in moyer2018invariant , as presented in table˜2. The proposed method outperforms the other approaches both in terms of accuracy and fairness, showcasing the effectiveness of the novel penalties. Further comparison under sensitive attributes is presented in section˜A.5.
@@
Since balunovic2022fair achieves the best balance between fairness and accuracy, in fig.˜4 we compare our proposed method from section˜4 against all datasets used in their evaluation. These include four distinct datasets: Crime and Health, which contain continuous features that NFs can handle directly, and Compas and Adult, which include discrete features requiring separate handling by Normalizing Flows.
The plots demonstrate that our proposed method consistently outperforms or matches balunovic2022fair , with the added advantage of not requiring the sensitive attribute during evaluation. Additionally, in section˜A.5, we report the Pareto fronts obtained by varying the latent dimension size, highlighting the impact on performance.
@@
In order to have a better sense of what the CF approach presented in section˜4 is learning, we visualize the learned representation in fig.˜3.
@@ 7 Conclusions
In this work, we introduce a novel framework for learning fair specialized representations using CFD, addressing key limitations of existing adversarial and normalizing flow-based approaches. By leveraging CFs, our method ensures stable and efficient fairness by minimizing sensitive information leakage while maintaining high predictive performance. We also present a simplified version of the framework that allows for certifiable bounds on the amount of sensitive information used by downstream tasks. Experimental results show that our approaches consistently outperform or match state-of-the-art methods in terms of the fairness-accuracy trade-off across a range of benchmark datasets. Unlike many existing approaches, our method does not require access to sensitive attributes at inference time, making it more practical for real-world applications. Our approach opens a new avenue for research in fair representation learning by offering a principled, non-adversarial alternative to existing methods. However, despite its simplicity and effectiveness, the proposed framework has certain limitations. Similar to many fairness-oriented algorithms, it encounters challenges in high-dimensional settings, where the CFD tends to lose effectiveness. Moreover, the moment-based approach is inherently tailored to classification tasks, leaving opportunities for extending such approaches to regression contexts. These limitations underscore the need for further research to improve the scalability and robustness of our method in such scenarios. Additionally, current fairness benchmarks largely focus on low-dimensional datasets, limiting the ability to comprehensively evaluate and compare methods in more realistic and complex settings. Developing more sophisticated, high-dimensional fairness datasets would greatly enhance the evaluation of approaches like ours and drive further innovation in the field.
Broader Impact
@@
Acknowledgements
We thank the anonymous reviewers for their insightful comments and suggestions, which helped improve the clarity and quality of this work. We additionally thank Alessandro Fabris and Davide Dalle Pezze, for the valuable early-stage feedback and discussions on this paper.